# Deliberate Planning of 3D Bin Packing on Packing Configuration Trees

### v3

@@ C.2 Understanding of Model Behaviors
Figure 21: (a)$\sim$(d) Different packing stages of the same sequence. The learned policies assign little items (colored in blue) to gaps and save room for future uncertainty. (e) Online 3D-BPP where load balancing is considered. (f) Online 3D-BPP with isle-friendliness, and different color means different item categories.
C.3 Performance on ICRA Stacking Challenge
The IEEE International Conference on Robotics and Automation (ICRA) organized the Stacking Challenge in 2023 (Sim4Dexterity, 2023). This competition primarily focuses on packing problem variants with $u>0$, including online packing problems with $p=s=1$, forward-looking packing problems with $p>s=1$, and general packing problems with $p>s>1$. The bin dimensions are set to $S_{x}=2.141,S_{y}=1.076,S_{z}=0.99$.
As shown in Figure 22, the item dimensions follow normal distributions: length $s_{x}\sim\mathcal{N}(0.45,0.09^{2})$, width $s_{y}\sim\mathcal{N}(0.3,0.05^{2})$, and height $s_{z}\sim\mathcal{N}(0.17,0.03^{2})$. Since the details and code of the participating algorithms have not been publicly released by the organizers, we compare our method with continuous-domain packing algorithms introduced in Section 4.3. The test results are summarized in Table 14. In the continuous domain, our method continues to effectively utilize the operational properties of selectable and previewed items for efficient planning, with performance improving as decision variable available increases.
Figure 22: Item dimensions of ICRA stacking challenge.
@@ C.3 Performance on ICRA Stacking Challenge
[…]
@@ C.4 Running Costs
Table 15: Running costs ($seconds$) tested on online 3D-BPP with discrete solution space (Section 4.1) and continuous solution space (Section 4.3). The running costs of the latter are usually more expensive since checking Constraints 2 and 3 in the continuous domain is more time-consuming.
[…]
@@
Heuristic $Random$ $4.59\times 10^{-2}$- $2.03\times 10^{-2}$- $4.62\times 10^{-2}$-
@@
C.5 Scalability
@@
Item sizes Discrete Continuous
[…]
$[S^{d}/10,S^{d}/2]$ 33.0 51.5 $86.0\%$ $9.5\times 10^{-2}$ 1.6 27.0 197.5 $66.3\%$ $2.4\times 10^{-2}$ 7.3
$[S^{d}/10,S^{d}/5]$ 241.3 67.2 $81.3\%$ $9.8\times 10^{-3}$ 0.3 185.4 956.5 $61.9\%$ $3.7\times 10^{-2}$ 5.2
[…]
Appendix D Visualized Results
We visualize the experimental results of different BPP variations on three settings in Figure 23, Figure 24, and Figure 25. It is clearly observed that as the number of packing decision variables increases, the packing results become more compact. We also provide the visualized results of large-scale packing in Figure 26. Each plot is about result tested by a randomly generated item sequence. The real-world packing results conducted via industrial production are provided in Figure 27.

### v4

@@ D.2 Understanding of Model Behaviors
[…]
D.3 Performance on ICRA Stacking Challenge
The IEEE International Conference on Robotics and Automation (ICRA) organized the Stacking Challenge in 2023 (Sim4Dexterity, 2023). This competition primarily focuses on packing problem variants with $u>0$, including online packing problems with $p=s=1$, forward-looking packing problems with $p>s=1$, and general packing problems with $p>s>1$. The bin dimensions are set to $S_{x}=2.141,S_{y}=1.076,S_{z}=0.99$. The item dimensions follow normal distributions: length $s_{x}\sim\mathcal{N}(0.45,0.09^{2})$, width $s_{y}\sim\mathcal{N}(0.3,0.05^{2})$, and height $s_{z}\sim\mathcal{N}(0.17,0.03^{2})$. Since the details and code of the participating algorithms have not been publicly released by the organizers, we compare our method with continuous-domain packing algorithms introduced in Section 4.3. The test results are summarized in Table 15. In the continuous domain, our method continues to effectively utilize the operational properties of selectable and previewed items for efficient planning, with performance improving as decision variables increases.
@@ D.3 Performance on ICRA Stacking Challenge
Method Prev. Sel. Un. Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@ D.4 Running Costs
Table 16: Running costs (secondsseconds) tested on online 3D-BPP with discrete solution space (Section 4.1) and continuous solution space (Section 4.3). The running costs of the latter are usually more expensive since checking Constraints (2) and (3) in the continuous domain is more time-consuming.
Method Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
@@
Heuristic RandomRandom $4.59\times 10^{-2}$ - $2.03\times 10^{-2}$ - $4.62\times 10^{-2}$ -
@@
D.5 Scalability
The number of PCT nodes changes constantly with the generation and removal of leaf nodes during the packing process. To verify whether our method can solve packing problems with a larger scale $|\textbf{B}|$, we conduct a stress test on setting 2setting\,2 where the most orientations are allowed and the most leaf nodes are generated. We limit the maximum item sizes $s^{d}$ to $S^{d}/5$ so that more items can be accommodated. We transfer the best-performing policies on setting 2setting\,2 (trained with EMS) to these new datasets without any fine-tuning. The results are summarized in Table 17.
Table 17: Scalability on larger packing problems. $|\textbf{L}|$ is the average number of leaf nodes per step. $|\textbf{L}|$ will not increase exponentially with $|\textbf{B}|$ since invalid leaf nodes will be removed.
Item sizes Discrete Continuous
$|\textbf{B}|$ $|\textbf{L}|$ Uti. Time $|\textbf{L}|/|\textbf{B}|$ $|\textbf{B}|$ $|\textbf{L}|$ Uti. Time $|\textbf{L}|/|\textbf{B}|$
$S^{d}/10,S^{d}/2]$ 33.0 51.5 $86.0\%$ $9.5\times 10^{-2}$ 1.6 27.0 197.5 $66.3\%$ $2.4\times 10^{-2}$ 7.3
$S^{d}/10,S^{d}/5]$ 241.3 67.2 $81.3\%$ $9.8\times 10^{-3}$ 0.3 185.4 956.5 $61.9\%$ $3.7\times 10^{-2}$ 5.2
PCT size will not grow exponentially with packing scale $|\textbf{B}|$ since invalid leaf nodes will be removed from leaf nodes L during the packing process, both discrete and continuous cases. For continuous cases, $|\textbf{L}|$ is more sensitive to $|\textbf{B}|$ due to the diversity of item sizes (i.e. $|\mathcal{I}|=\infty$), however, $|\textbf{L}|$ still doesn’t explode with $|\textbf{B}|$ and it grows in a sub-linear way. Our method can execute packing decisions at a real-time speed with controllable PCT sizes, even if the item scale is around two hundred.
D.6 Computational Cost of Attention
PCT employs attention mechanisms to flexibly encode varying state and action spaces, including Graph Attention Network (GAT) (Velickovic et al., 2018) and a pointer mechanism (Vinyals et al., 2015). GAT incurs a computational complexity of $O(N^{2})$, while the pointer mechanism scales as $O(|\textbf{L}_{t}|)$. Here $N$ is the total number of nodes in the PCT, and $|\textbf{L}_{t}|$ is the number of leaf nodes at time step $t$. We present quantitative measurements to illustrate how their computational costs grow with increasing node count, as visualized in Figure 25.
During training, the policy launches $k_{p}$ parallel packing environments to collect experience and performs a forward rollout of $k_{s}$ steps per update. We recommend $k_{p}=64$ and $k_{s}=5$ as a practical setting. Under this batch size, both the inference time and memory usage of GAT scale quadratically with the number of nodes, whereas the pointer mechanism scales linearly, as shown in Figure 25 (a) and (b). The quadratic cost of GAT also explains the difficulty of directly training packing policies at large scales. During testing, i.e., $k_{p}=k_{s}=1$, the running time of both GAT and pointer mechanisms remains stable as the number of nodes increases, as shown in Figure 25 (c). This is due to their lightweight design and low CPU memory footprint, as further demonstrated in Figure 25 (d), which helps prevent GPU compute unit saturation.
(a) Inference time for training (b) Memory cost for training(c) Inference time for test (d) Memory cost for test
Figure 25: Computational cost of attention mechanisms. The reported node count refers to the total node number in PCT for GAT, and to the leaf node number for the pointer mechanism.
D.7 Sensitivity of Constraint Reward Weights
We introduce constraint rewards ${f}(\cdot)$ into the packing reward function to handle additional complex constraints, as described in Section 4.4, forming a multi-objective optimization problem. The item weight is revised as:
@@ D.7 Sensitivity of Constraint Reward Weights
Table 18: Parameter sensitivity study of the constraint reward weight $c$.
Constraints Weight $c$ Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
Obj. Uti. $\uparrow$ Num. $\uparrow$ Obj. Uti. $\uparrow$ Num. $\uparrow$ Obj. Uti. $\uparrow$ Num. $\uparrow$
Isle Friendliness $\downarrow$ 0.1 0.14 71.8% 28.7 0.08 85.0% 32.6 0.14 74.2% 28.6
1.0 0.13 66.2% 26.4 0.07 75.4% 29.0 0.13 68.5% 26.4
10.0 0.10 59.5% 23.9 0.05 63.1% 24.1 0.09 62.7% 24.0
Load Balancing $\downarrow$ 0.1 1.38 70.9% 27.5 0.68 83.3% 32.2 0.22 71.1% 27.7
1.0 1.16 65.2% 25.4 0.55 75.6% 29.2 0.18 64.5% 25.2
10.0 0.90 59.0% 22.9 0.49 68.2% 26.4 0.14 58.5% 22.8
Height Uniformity $\downarrow$ 0.1 3.74 73.8% 28.5 1.99 83.3% 32.2 3.77 72.8% 28.2
1.0 2.57 73.2% 28.1 1.36 82.1% 31.8 2.54 71.9% 27.9
10.0 2.18 64.5% 25.0 1.19 72.5% 28.1 2.24 65.4% 25.3
Kinematic Constraints $\uparrow$ 0.1 0.95 72.7% 28.2 0.96 84.7% 32.5 0.94 74.4% 28.7
1.0 1.03 70.2% 27.2 1.06 81.3% 31.3 1.03 71.3% 27.6
10.0 1.12 68.5% 26.6 1.17 76.2% 29.3 1.11 69.2% 26.7
Load Bearing Constraints $\downarrow$ 0.1 1.40 69.6% 27.8 1.36 80.3% 31.6 0.78 68.4% 27.2
1.0 1.22 61.9% 24.7 1.25 74.1% 29.0 0.72 63.1% 25.2
10.0 1.13 57.2% 23.0 1.03 62.2% 24.5 0.66 58.5% 23.2
Bridging Constraints $\uparrow$ 0.1 1.19 68.8% 26.8 1.30 80.5% 31.3 1.19 68.8% 26.8
1.0 1.23 66.2% 25.7 1.33 78.3% 30.4 1.25 65.6% 25.6
10.0 1.31 62.0% 24.1 1.42 70.1% 27.0 1.31 61.5% 24.0
We can observe that increasing the weight of constraint rewards inevitably reduces space utilization. Some constraints conflict mildly with this objective. For example, Kinematic Constraint aims to minimize the impact of placed items on subsequent robot motions by favoring horizontally distal placements for early items. Optimizing this placement order has a relatively minor impact on space utilization, which reaches the highest utilization at $c=10.0$. In contrast, for strongly conflicting constraints like Load-bearing Constraint, where minimizing per-object load can be achieved by placing fewer items, increasing $c$ leads to a more significant utilization decline, i.e., the lowest utilization at $c=10.0$.
Appendix E Additional Real-World Specifications
We adopt a modular gripper that can actively adjust its shape to avoid robot-object collisions. The gripper comprises five grasping modules of varying sizes, capable of reconfiguring themselves based on the dimensions of the target box, as shown in Figure 26 (a). Each module is equipped with multiple suction cups, totaling 165 units (each with a diameter of 39 mm), providing a maximum suction force of $260kg/m^{2}$. The central module is fixed in place and supports the vertical motion of the four surrounding modules, as demonstrated in Figure 26 (b). These movable modules are actuated by rodless pneumatic cylinders, allowing them to be raised or lowered as needed. The cylinders are powered by vacuum pumps, which are also shared to generate the suction force for gripping items, thereby eliminating the need for an additional power source for gripper adjustment.
To further prevent collisions in real-world factories, the robot avoids moving directly above the target position to perform a vertical drop. Instead, it follows a simplified yet effective placement manner by planning an oblique insertion point as an intermediate waypoint along the motion trajectory, as illustrated in Figure 27. This oblique approach reduces the risk of collisions during vertical descent caused by interference from adjacent boxes. To determine valid oblique insertion directions, we examine eight candidates $d\in\{(c_{1},c_{2},0)\}$ with $c_{1},c_{2}\in\{-1,0,1\}$, subject to the constraint $c_{1}\neq 0$ or $c_{2}\neq 0$. Directions without obstruction from taller boxes are considered valid, and we prioritize those with non-zero values in both axes (i.e., $c_{1}\cdot c_{2}\neq 0$) to maximize spatial clearance.
Figure 26: The modular designed gripper for constraint placement in packing scenarios. This flexible gripper consists of five suction cup modules (a) which can move vertically and online adjust its shape (b) based on the target item size. The central module 1 is fixed and serves as a guide, supporting the vertical movement of the surrounding modules.
Figure 27: Simplified robot packing trajectory. The green box is to be placed into the target placement outlined by the green bounding box. The yellow circle marks the waypoint for oblique insertion. The robot first moves to this waypoint, then proceeds to the final placement.
Appendix F Visualized Results
We visualize the experimental results of different BPP variations on three settings in Figure 28, Figure 29, and Figure 30. It is clearly observed that as the number of packing decision variables increases, the packing results become more compact. We also provide the visualized results of large-scale packing in Figure 31. Each plot is about results tested by a randomly generated item sequence. The real-world packing results conducted via industrial production are provided in Figure 32.